# Scaling Open-Source Multimodal Data by 10x: A Multimodal Dataset with One Trillion Tokens

### v1

@@ Abstract
Multimodal interleaved datasets featuring free-form interleaved sequences of images and text are crucial for training frontier large multimodal models (LMMs). Despite the rapid progression of open-source LMMs, there remains a pronounced scarcity of large-scale, diverse open-source multimodal interleaved datasets. In response, we introduce [Uncaptioned image] MINT-1T, the most extensive and diverse open-source Multimodal INTerleaved dataset to date. [Uncaptioned image] MINT-1T comprises of one trillion text tokens and three billion images, a 10x scale-up from existing open-source datasets. Additionally, we include previously untapped sources such as PDFs and ArXiv papers. As scaling multimodal interleaved datasets requires substantial engineering effort, sharing the data curation process and releasing the dataset greatly benefits the community. Our experiments show that LMMs trained on MINT-1T rival the performance of models trained on the previous leading dataset, OBELICS. Our data and code will be released at https://github.com/mlfoundations/MINT-1T.
@@
Figure 2: Samples of multimodal document from the HTML, PDF, and ArXiv subsets of Refer to caption MINT-1T with each document containing a sequence of images interleaved with text. Previous work has shown that interleaved data improves question-answering performance in the context of Flamingo-style models [Laurençon et al., 2023] and for training large multimodal models with a strong performance on both text-only and multimodal benchmarks [McKinzie et al., 2024]. MINT-1T is the first open-source work to scale interleaved datasets to one trillion text tokens and collect interleaved documents from PDFs and ArXiv at large scales. Samples in this figure are text truncated due to space.
@@ 1 Introduction
Model Experiments Our experiments demonstrate that LMMs trained on MINT-1T not only match but potentially surpass the performance of models trained on the best existing open-source dataset, OBELICS, while offering a tenfold increase in scale.
@@ 2 Dataset Construction
[Uncaptioned image] MINT-1T 1T 3.4B 927M HTML/PDFs/ArXiv
Table 1: Survey of multimodal interleaved datasets: Current open-source interleaved datasets are smaller than proprietary datasets. Beyond HTML data sources, MINT-1T uniquely includes data from PDFs and ArXiv documents.
Figure 3: The filtering process for MINT-1T shows how tokens are removed throughout the data pipeline for HTML, PDFs, and ArXiv papers.
MINT-1T curates a large-scale open-source dataset that taps into more diverse sources of interleaved documents such as PDFs and ArXiv papers. This section outlines our methods for sourcing multimodal documents, filtering low-quality documents, deduplicating data, and removing not safe for work and undesirable content. The final dataset contains 965 billion (B) HTML document tokens, 51B PDF tokens, and 10B ArXiv tokens (for more detailed numbers and filtering ratios refer to Figure 3).
@@ 2.1.1 HTML Pipeline
We follow OBELICS’s method for extracting interleaved multimodal documents from CommonCrawl WARC files by parsing each WARC entry’s DOM tree. While OBELICS only processed documents from February 2020 to February 2023 CommonCrawl dumps, we have expanded our document pool to include HTML documents from May 2013 to December 2023 (note that we use the full dumps from May 2017 to December 2023 and partial dumps from earlier years). Following OBELICS, we filter out any documents that contain no images or more than thirty images or any images with URLs that include inappropriate substrings such as logo, avatar, porn, and xxx.
2.1.2 PDF & ArXiv Pipeline
We source PDF documents from CommonCrawl WAT files from February 2023 to April 2024 dumps. Initially, we extract all the PDF links from these dumps and perform URL deduplication. We then attempt to download and read PDFs using PyMuPDF  [^1] https://github.com/pymupdf/PyMuPDF-Utilities/blob/master/text-extraction/multi_column.py. We discard PDF documents that are more than 50MB large (as they likely contain predominantly large images) and PDFs that are over 50 pages long. We exclude pages that contain no text and extract a reading order for the remaining pages. Reading order is obtained by finding the bounding box of all text blocks on a page, clustering the blocks based on columns, and ordering the blocks from top left to bottom right. Images are anchored in the sequence based on the proximity between the image’s bounding box and text blocks on the same page. We discuss the limitations of this approach in Appendix A.1. We apply this same pipeline to ArXiv PDFs. However, given that ArXiv is already a highly curated data source, we do not perform any of the filtering and deduplication described in the rest of this section.
@@ 2.5 Deduplication
Our text deduplication pipeline consists of two stages: paragraph and document deduplication within each Common Crawl snapshot, followed by a global document text deduplication across all the snapshots. We also perform image deduplication to remove repetitive, uninformative images such as icons and logos. All deduplication steps are done separately for each data source.
@@ 2.5.3 Image Deduplication
2.6 Estimating Dataset Statistics
Due to time and resource constraints, not all the images in the HTML subset of our dataset are downloaded. We estimate the effects of image filtering and deduplication by processing one to two thousand shards per CommonCrawl dump out of roughly 70 to 90 thousand shards total. We empirically found that this is sufficient to estimate the full dump size. For a more precise removal rate post-image deduplication, we fully processed six dumps and found that, on average, 30% of the tokens are removed. We apply this rate to the rest of the dumps.
@@ 4.1 Training Setup
Modeling Our architecture is most similar to Idefics2 [Laurenccon et al., 2024]. We use the ViT-H vision encoder with resolution 378 from Data Filtering Networks [Fang et al., 2023] and pass the patch embeddings into a six layer perceiver resampler [Alayrac et al., 2022]. Each image is represented as 128 tokens. The pooled patch embeddings are interleaved with the text token embeddings and passed into Phi-3 mini language model [Abdin et al., 2024]. We keep the vision encoder frozen while training the resampler and the language model. We use a batch size of, on average, 1.8M multimodal tokens. For all of our training runs, we use 2,000 warmup steps with a maximum learning rate of $5*10^{-5}$ and cosine learning rate decay. We also apply 0.05 weight decay to all trainable parameters. All of our training is done using the OpenFlamingo [Awadalla et al., 2023] codebase. We train all of our models on 32 H100 GPUs for a total of 1,920 GPU hours per experiment.
@@ 4.2 Evaluation Setup
We assess a model’s capability to reason about multimodal interleaved sequences through its performance on in-context learning abilities and multi-image reasoning performance.
(a) We plot the average performance on COCO, TextCaps, OK-VQA, VizWiz, and TextVQA when using one to fifteen demonstrations. We find that the model trained on HTML and PDF documents from MINT-1T performs better on average across all demonstrations in comparison to models trained on the HTML portion of MINT-1T or OBELICS.
(b) MINT-1T outperforms OBELICS on many domains. We improve on OBELICS in Science and Technology domains as the PDF subset of MINT-1T contains a large representation of these domains. MINT-1T also performs better than OBELICS on Art domains despite not having a significant representation of these domains.
In-context Learning The models are evaluated on four-shot and eight-shot in-context learning performance on various captioning benchmarks (COCO (Karpathy test) [Lin et al., 2014] and TextCaps (validation) [Sidorov et al., 2020]) and visual question answering datasets (VQAv2 (validation) [Agrawal et al., 2015], OK-VQA (validation) [Marino et al., 2019], TextVQA (validation) [Singh et al., 2019], and VizWiz (validation) [Gurari et al., 2018]). For all evaluations, we randomly sample demonstrations from the training set. Our reported scores are averaged over three evaluation runs where we randomize demonstrations. We find that performance is sensitive to the chosen prompts, so for models trained on MINT-1T and OBELICS, we ablate through different prompts for each task and choose the prompt that performs best. The list of prompts we use and generation parameters can be found in Appendix A.2.
@@ 4.3 Experiments
Model Shots Datasets
@@
OBELICS 4 $\textbf{108.1}\pm 0.91$ $\underline{80.4}\pm 0.96$ $\underline{48.4}\pm 0.79$ $\underline{42.4}\pm 0.74$ $61.8\pm 0.12$ $26.0\pm 0.28$
8 $\underline{109.4}\pm 0.71$ $83.9\pm 0.36$ $\textbf{49.6}\pm 0.08$ $\underline{43.8}\pm 0.43$ $62.3\pm 0.42$ $30.9\pm 1.10$
MINT-1T (HTML) 4 $105.2\pm 1.84$ $77.9\pm 1.27$ $48.1\pm 0.11$ $\underline{42.2}\pm 0.49$ $\textbf{63.5}\pm 0.00$ $25.5\pm 0.70$
8 $108.1\pm 0.70$ $81.4\pm 1.27$ $49.5\pm 0.01$ $\underline{43.4}\pm 0.19$ $\underline{64.5}\pm 0.7$ $\underline{30.9}\pm 1.37$
[Uncaptioned image] MINT-1T 4 $106.6\pm 0.43$ $\underline{81.1}\pm 1.05$ $\underline{48.4}\pm 0.06$ $\underline{42.7}\pm 0.28$ $62.8\pm 0.11$ $\textbf{27.3}\pm 0.60$
8 $\underline{109.3}\pm 0.40$ $\textbf{84.4}\pm 0.11$ $49.2\pm 0.17$ $43.2\pm 0.44$ $\underline{63.9}\pm 0.87$ $\underline{32.2}\pm 0.96$
@@
We first evaluate how the HTML portion of MINT-1T compares to OBELICS, given that we used different filtering and deduplication methods. We train two models on the HTML portions of MINT-1T and OBELICS for 10B multimodal tokens total and assess their in-context learning performance. In Table 2, we present 4-shot and 8-shot performance on common benchmarks; the model trained on our HTML documents performs overall comparably with many cases slightly outperforms OBELICS.
@@
[Uncaptioned image] MINT-1T 41.3 47.5
Table 3: Performance on multi-image reasoning benchmarks MMMU and Mantis-Eval.
Adding PDF and ArXiv documents Subsequently, we train on MINT-1T’s full data sources, with a mixture of HTML, PDF, and ArXiv documents. We sample $50\%$ of our interleaved documents from HTML, $45\%$ from PDFs, and $5\%$ from ArXiv. We train for a total of 10B multimodal tokens. As seen in Table 2, the model trained on the full MINT-1T data mixture performs on par with OBELICS and MINT-1T (HTML) on many in-context learning benchmarks (with the exception of TextCaps where the MINT-1T model outperforms other models). On more complex multimodal reasoning benchmarks, we find that the MINT-1T model outperforms OBELICS on MMMU and Mantis-Eval.
@@ 4.4 Fine-grained Trends
How Does In-Context Learning Performance Scale with Demonstrations? We evaluate models’ in-context learning performance when prompted with one to fifteen demonstrations. We run a single trial per shot count for each evaluation benchmark described in Section 4.2 with the exception of VQAv2, as the validation set is quite large. As seen in Figure 7(a), we find that the model trained on MINT-1T outperforms the model trained on the HTML subset of MINT-1T for all shots. Moreover, we find that the MINT-1T model performs similarly to OBELICS from one to ten shots but drastically outperforms it past ten shots. This is intuitive as MINT-1T contains PDF and ArXiv documents, which, on average, have more images.
Performance on Different Domains A major motivation for including documents from diverse domains in MINT-1T is to improve model generalization. In Figure 7(b), we break down performance on MMMU for each domain. With the exception of the Business domain, MINT-1T outperforms OBELICS and MINT-1T (HTML baseline). We highlight the performance increase on Science and Technology domains for MINT-1T and speculate that this can be attributed to the prevalence of these domains in ArXiv and PDF documents.

### v2

@@ Abstract
Multimodal interleaved datasets featuring free-form interleaved sequences of images and text are crucial for training frontier large multimodal models (LMMs). Despite the rapid progression of open-source LMMs, there remains a pronounced scarcity of large-scale, open-source multimodal interleaved datasets. In response, we introduce [Uncaptioned image] MINT-1T, the most extensive and diverse open-source Multimodal INTerleaved dataset to date. [Uncaptioned image] MINT-1T comprises of one trillion text tokens and 3.4 billion images, a 10x scale-up from existing open-source datasets. Additionally, we include previously untapped sources such as PDFs and ArXiv papers. As scaling multimodal interleaved datasets requires substantial engineering effort, sharing the data curation process and releasing the dataset greatly benefits the community. Our experiments show that LMMs trained on MINT-1T rival the performance of models trained on the previous leading dataset, OBELICS. We release our data at https://github.com/mlfoundations/MINT-1T.
@@
Figure 2: Samples of multimodal document from the HTML (Left), PDF (Middle), and ArXiv (Right) subsets of Refer to caption MINT-1T with each document containing a sequence of images interleaved with text. Previous work has shown that interleaved data improves question-answering performance in the context of Flamingo-style models [Laurençon et al., 2023] and for training large multimodal models with a strong performance on both text-only and multimodal benchmarks [McKinzie et al., 2024]. MINT-1T is the first open-source work to scale interleaved datasets to one trillion text tokens and collect interleaved documents from PDFs and ArXiv at large scales. Samples in this figure are text truncated due to space.
@@ 1 Introduction
Model Experiments Our experiments demonstrate that LMMs trained on [Uncaptioned image] MINT-1T not only match but potentially surpass the performance of models trained on the best existing open-source dataset, OBELICS, while offering a tenfold increase in scale.
@@ 2 Dataset Construction
[Uncaptioned image] MINT-1T 1.03T 3.45B 1057M HTML/PDFs/ArXiv
Table 1: Survey of multimodal interleaved datasets: Current open-source interleaved datasets are smaller than proprietary datasets. Beyond HTML data sources, [Uncaptioned image] MINT-1T uniquely includes data from PDFs and ArXiv documents.
Figure 3: The filtering process for Refer to caption MINT-1T shows how tokens are removed throughout the data pipeline for HTML, PDFs, and ArXiv papers.
MINT-1T curates a large-scale open-source dataset that taps into more diverse sources of interleaved documents such as PDFs and ArXiv papers. This section outlines our methods for sourcing multimodal documents, filtering low-quality documents, deduplicating data, and removing not safe for work and undesirable content. The final dataset contains 922 billion (B) HTML tokens, 106B PDF tokens, and 9B ArXiv tokens (for more detailed numbers and filtering ratios refer to Figure 3).
@@ 2.1.1 HTML Pipeline
We follow OBELICS’s method for extracting interleaved multimodal documents from CommonCrawl WARC files by parsing each WARC entry’s DOM tree. While OBELICS only processed documents from February 2020 to February 2023 CommonCrawl dumps, we have expanded our document pool to include HTML documents from May 2017 to April 2024 (note that we use the full dumps from October 2018 to April 2024 and partial dumps from earlier years). Following OBELICS, we filter out any documents that contain no images or more than thirty images or any images with URLs that include inappropriate substrings such as logo, avatar, porn, and xxx.
2.1.2 PDF Pipeline
We source PDF documents from CommonCrawl WAT files from February 2023 to April 2024 dumps. Initially, we extract all the PDF links from these dumps. We then attempt to download and read PDFs using PyMuPDF  [^1] https://github.com/pymupdf/PyMuPDF-Utilities/blob/master/text-extraction/multi_column.py. We discard PDF documents that are more than 50MB large (as they likely contain predominantly large images) and PDFs that are over 50 pages long. We exclude pages that contain no text and extract a reading order for the remaining pages. Reading order is obtained by finding the bounding box of all text blocks on a page, clustering the blocks based on columns, and ordering the blocks from top left to bottom right. Images are anchored in the sequence based on the proximity between the image’s bounding box and text blocks on the same page. We discuss the limitations of this approach in Appendix A.1.
2.1.3 ArXiv pipeline
ArXiv interleaved documents are built from the LaTeX source code. We use TexSoup  [^2] https://github.com/alvinwan/TexSoup to find figure tags and interleave the images with the paper text. For multi-file papers (i.e. where each section is written in a different Tex file), we identify the main Tex file and replace input tags with the contents of its file. We additionally clean up the the LaTex code removing imports, bibliography, tables, and citation tags. As ArXiv is already a highly curated data source, we do not perform any of the filtering and deduplication described in the rest of this section.
@@ 2.5 Deduplication
To remove duplicated content, we perform paragraph and document text deduplication within each CommonCrawl snapshot and image deduplication to remove repetitive, uninformative images such as icons and logos. All deduplication steps are done separately for each data source.
@@ 4.1 Training Setup
Modeling Our architecture is adopted from XGen-MM [Salesforce, 2024]. We use the ViT-H vision encoder with resolution 378 from Data Filtering Networks [Fang et al., 2023] and pass the patch embeddings into a six layer perceiver resampler [Alayrac et al., 2022]. Each image is represented as 128 tokens. The pooled patch embeddings are interleaved with the text token embeddings and passed into Phi-3 mini language model [Abdin et al., 2024]. We keep the vision encoder frozen while training the resampler and the language model. We use a batch size of, on average, 1.8M multimodal tokens. For all of our training runs, we use 2,000 warmup steps with a maximum learning rate of $5*10^{-5}$ and cosine learning rate decay. We also apply 0.05 weight decay to all trainable parameters. All of our training is done using the OpenFlamingo [Awadalla et al., 2023] codebase. We train all of our models on 32 H100 GPUs for a total of 1,920 GPU hours per experiment.
@@ 4.2 Evaluation Setup
We assess a model’s capability to reason about multimodal interleaved sequences through its in-context learning abilities and multi-image reasoning performance.
(a) Average performance on captioning and visual question answering benchmarks when using one to eight demonstrations. The model trained on MINT-1T performs better across all demonstrations in comparison to the HTML portion of MINT-1T or OBELICS.
(b) MINT-1T outperforms OBELICS on many domains. We improve on OBELICS in Science and Technology domains as the PDF subset of MINT-1T contains a large representation of these domains. MINT-1T also performs better than OBELICS on Art.
In-context Learning The models are evaluated on four-shot and eight-shot in-context learning performance on various captioning benchmarks (COCO (Karpathy test) [Lin et al., 2014] and TextCaps (validation) [Sidorov et al., 2020]) and visual question answering datasets (VQAv2 (validation) [Agrawal et al., 2015], OK-VQA (validation) [Marino et al., 2019], TextVQA (validation) [Singh et al., 2019], and VizWiz (validation) [Gurari et al., 2018]). For all evaluations, we randomly sample demonstrations from the training set. Our reported scores are averaged over multiple evaluation runs where we randomize demonstrations. We find that performance is sensitive to the chosen prompts, so we ablate through different prompts for each task and choose the prompt that performs best. The list of prompts we use and generation parameters can be found in Appendix A.2.
@@ 4.3 Experiments
Model Shots Datasets Average
@@
OBELICS 4 $\textbf{108.1}\pm 0.91$ $\underline{80.4}\pm 0.96$ $48.4\pm 0.79$ $42.4\pm 0.74$ $61.8\pm 0.12$ $26.0\pm 0.28$ $61.2\pm 0.28$
8 $\underline{109.4}\pm 0.71$ $\underline{83.9}\pm 0.36$ $49.6\pm 0.08$ $43.8\pm 0.43$ $62.3\pm 0.42$ $\underline{30.9}\pm 1.10$ $63.3\pm 0.24$
MINT-1T (HTML) 4 $105.4\pm 0.99$ $\underline{79.2}\pm 1.17$ $48.0\pm 0.12$ $43.7\pm 0.35$ $63.5\pm 0.11$ $25.6\pm 0.44$ $60.9\pm 0.27$
8 $107.9\pm 0.58$ $81.9\pm 1.29$ $49.7\pm 0.37$ $44.2\pm 0.30$ $64.3\pm 0.15$ $30.3\pm 0.87$ $63.0\pm 0.28$
[Uncaptioned image] MINT-1T 4 $107.0\pm 0.13$ $\underline{79.7}\pm 0.62$ $\textbf{49.7}\pm 0.19$ $\textbf{45.0}\pm 0.39$ $\textbf{64.9}\pm 0.07$ $\textbf{27.5}\pm 0.32$ $\textbf{62.3}\pm 0.13$
8 $\underline{108.8}\pm 0.34$ $\underline{84.3}\pm 0.51$ $\textbf{51.1}\pm 0.18$ $\textbf{45.6}\pm 0.10$ $\textbf{66.1}\pm 0.09$ $\underline{31.8}\pm 0.53$ $\textbf{64.6}\pm 0.14$
@@
We first evaluate how the HTML portion of MINT-1T compares to OBELICS, as OBELICS is the previous leading interleaved dataset and is also curated from HTML documents. We train two models on the HTML portions of MINT-1T and OBELICS for 10B multimodal tokens total and assess their in-context learning performance. In Table 2, we present 4-shot and 8-shot performance on common benchmarks; the model trained on MINT-1T HTML documents performs better than OBELICS on VQA tasks but worse on captioning benchmarks. On average OBELICS performs slightly better than MINT-1T (HTML). We explore how model architecture impacts this result in Section 4.5.
@@
[Uncaptioned image] MINT-1T 41.3 43.3
Table 3: Performance on image reasoning benchmarks MMMU and Mantis-Eval.
Adding PDF and ArXiv documents Subsequently, we train on MINT-1T’s full data sources, with a mixture of HTML, PDF, and ArXiv documents. We sample $50\%$ of our interleaved documents from HTML, $45\%$ from PDFs, and $5\%$ from ArXiv. We train for a total of 10B multimodal tokens. As seen in Table 2, the model trained on the full MINT-1T data mixture outperforms OBELICS and MINT-1T (HTML) on most in-context learning benchmarks. On more complex multimodal reasoning benchmarks, the MINT-1T model outperforms OBELICS on MMMU but performs worse on Mantis-Eval.
@@ 4.4 Fine-grained Trends
How Does In-Context Learning Performance Scale with Demonstrations? We evaluate models’ in-context learning performance when prompted with one to eight demonstrations. We run a single trial per shot count for each evaluation benchmark described in Section 4.2. As seen in Figure 7(a), we find that the model trained on MINT-1T outperforms the model trained on the HTML subset of MINT-1T and OBELICS on all shots. Moreover, we find that the MINT-1T (HTML) model performs slightly worse than OBELICS.
Figure 8: Performance on captioning and visual question answering (VQA) tasks. The HTML subset of MINT-1T outperforms OBELICS on VQA tasks but performs worse on captioning benchmarks.
Performance on Captioning and Visual Question Answering Tasks In Figure 8, we present the average in-context learning performance on captioning and visual question answering (VQA) benchmarks. OBELICS outperforms all MINT-1T variants on four shot captioning benchmarks and performs slightly worse to MINT-1T on eight shot captioning. However, we find that on VQA benchmarks MINT-1T is significantly better than both baselines. We also see that MINT-1T (HTML) also outperforms OBELICS on VQA tasks.
Performance on Different Domains A motivation for including diverse domains in MINT-1T is to improve model generalization. In Figure 7(b), we break down performance on MMMU for each domain. With the exception of the Business domain, MINT-1T outperforms OBELICS and MINT-1T (HTML). We highlight the performance increase on Science and Technology domains for MINT-1T and speculate that this can be attributed to the prevalence of these domains in ArXiv and PDF documents.
Figure 9: Impact of architecture: On in-context learning benchmarks, XGen-MM models perform marginally better when trained on OBELICS compared to MINT-1T’s HTML subset. In contrast, Idefics2 models show a slight advantage for MINT-1T (HTML) over OBELICS.
Model Shots Datasets Average
COCO TextCaps OKVQA TextVQA VQAv2 VizWiz
OBELICS 4 $110.2\pm 0.38$ $83.1\pm 1.28$ $52.8\pm 0.04$ $46.7\pm 0.23$ $63.8\pm 0.10$ $28.9\pm 0.50$ $64.3\pm 0.24$
8 $111.8\pm 1.04$ $86.6\pm 0.31$ $\textbf{54.8}\pm 0.04$ $46.9\pm 0.02$ $64.1\pm 0.09$ $36.2\pm 0.51$ $66.7\pm 0.20$
MINT-1T (HTML) 4 $\textbf{110.9}\pm 0.01$ $\textbf{84.8}\pm 0.05$ $52.9\pm 0.26$ $\textbf{47.0}\pm 0.31$ $\textbf{65.6}\pm 0.04$ $29.0\pm 0.64$ $\textbf{65.0}\pm 0.12$
8 $111.3\pm 0.05$ $\textbf{87.5}\pm 0.11$ $54.1\pm 0.49$ $\textbf{48.1}\pm 0.12$ $64.8\pm 1.04$ $36.6\pm 0.13$ $\textbf{66.9}\pm 0.19$
Table 4: Idefics2 model results: We compare OBELICS and MINT-1T (HTML), when training an Idefics2 LMM. Models are evaluated using four and eight in-context learning examples, with each evaluation run for two trials. We report the mean performance and standard deviation.
4.5 Impact of Model Architecture
Our experiments, presented in Section 4.3, use XGen-MM’s architecture. Unlike with large language models, the design space for multimodal models is much more diverse with many architectures for aligning a vision encoder to a language model. Naturally, we were curious if our results would hold for other popular training setups.
To investigate this, we replicate our training experiments using Idefics2’s architecture. Idefics2 differs from XGen-MM in that it freezes a non-instruction finetuned large language model and adds LoRA [Laurenccon et al., 2024] matrices on all linear layers. For our Idefics2 reproduction we use the Mistral-7B-v0.3 language model and DFN ViT-H vision encoder with resolution 384. Unlike Idefics2, we do not experiment with flexible image resolution in training and keep the vision encoder completely frozen. We present in-context learning results for Idefics2 model experiments in Table 4. We find that MINT-1T’s HTML subset performs better than OBELICS with notable gains on TextVQA, VQAv2, and TextCaps. We highlight the performance gap difference between XGen-MM and Idefics2 ablations in Figure 9. One key difference in the Idefics2 experiments is that the HTML subset of MINT-1T performs much more competitively on captioning benchmarks in comparison to OBELICS.